OpenStudy (anonymous):

|-16| + 92 = 20y - |-32| - 19y BTW ^ THE 2 IS ON THE TOP

OpenStudy (anonymous):

absolute value function is defined as: x if x >= 0 -x if x < 0 |-16| < 0 , therefore it falls under the "-x" category. thus, |-16| + 92 = 20y - |-32| - 19y -(-16) + 92 = 20y - -(32) - 19y 16 + 92 = 20y + 32 - 19y 108 = y + 32 76 = y <- ans not so sure by what you mean of "the two is on top" but 76 seems to be the correct answer

OpenStudy (anonymous):

as in the 2 is on the right top corner of the 9